Two cars started from the same point. One traveled north at 55 mi/h and the other traveled west at 65 mi/h. How far apart were the cars after 4 hours? Round to the nearest tenth of a mile.
Found 2 solutions by mananth, Alan3354:
Answer by mananth(16946)

They are travelling at right angles from the starting point .
At any given time they form a right angle with the starting point.
Car 1 after 4 hours 4*55=220 miles
Car 2 after 4 hours 4*65 =260 miles
240^2+260^2=D^2
3025+4225= D^2
d^2=7250
d= 85.1 miles -their distance apart

Answer by Alan3354(69443)